# Gradient Extrapolation for Debiased Representation Learning

### v1

@@ 4.1.3 GERNE as general framework for debiasing
The proof is in Appendix A.
Based on Eq. 8, we can establish the link between GERNE and other methods for different values of $\beta$ and $c$ as illustrated in Fig. 1:
    For $\beta=-1$, $\mathcal{L}_{ext}=\mathcal{L}_{b}$ and GERNE is equivalent to balanced-classes ERM approach.
    For $c=1,\beta=0$, $p_{ext}(a|y)=\frac{1}{A}$; GERNE is equivalent to the resampling method [16] (upsampling the minority group).
    For $c\cdot(\beta+1)=1$, $p_{ext}(a|y)=\frac{1}{A}$; In this case, $\mathcal{L}_{ext}$ is equivalent, in expectation, to the loss of the resampling method ($\mathcal{L}_{rs}$ in Fig. 1). However, the variances of the losses differ. In fact, GERNE permits controlling the variance of its loss, thanks to its hyperparameters, which might help escaping sharp minima [1] and generalize better [23]. The derivation of the variance of GERNE’s loss is detailed in Appendix C.
    For $c\cdot(\beta+1)>1$, $p_{ext}(a|y)>\frac{1}{A}\text{ if }\alpha_{ya}<\frac{1}{A}$ (also $p_{ext}(a|y)<\frac{1}{A}\text{ if }\alpha_{ya}>\frac{1}{A}$). In this case, GERNE simulates batches where the underrepresented groups (i.e. $\alpha_{ya}<\frac{1}{A}$) in the dataset to be oversampled in these simulated batches.
@@ 4.1.4 Upper and lower bounds of $\beta$
Having $p_{ext}(a|y)$ in Eq. 8 in $[0,1]$, $\beta$ should satisfy:
$\begin{split}\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\min(i_{ya}^{1},i_{ya}^{2})\leq\beta\leq\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})\end{split}$ (9)
For: $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$.
These limits are used when tuning $\beta,c$.
4.1.5 Tuning $\beta$ to minimize the risk of worst-case group
Eq. 5 can be rewritten as follows:
$\mathcal{L}_{ext}=\frac{1}{K}\sum_{g=(y,a)\in\mathcal{G}}p_{ext}(a|y)(\beta)\cdot L_{g}$ (10)
@@ 4.1.5 Tuning $\beta$ to minimize the risk of worst-case group
In the presence of spurious correlations, minority or less-represented groups often experience higher risks, primarily due to the model’s limited exposure to these groups during training [20]. Taking this into consideration, we define $g^{\prime}=(y^{\prime},a^{\prime})=\arg\min_{(y,a)\in\mathcal{G}}\alpha_{ya}$. Since $L_{g^{\prime}}$ is multiplied by $p_{ext}(a^{\prime}|y^{\prime})$ defined in Eq. 8, increasing $\beta$ assigns more weight to $L_{g^{\prime}}$ in Eq. 10 than $L_{g}$ with $g\neq g^{\prime}$, thereby encouraging a reduction in $L_{g^{\prime}}$ during training. We outline the detailed steps of our approach for the known attribute case in Algorithm 1.
@@
Input: $\mathcal{X}_{y,a}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$ and $a\in\mathcal{A}$, $f$ with initial $\theta=\theta_{0}$, ##\## epochs $E$, batch size per label $B$, ##\## classes $K$, ##\## attributes $A$, learning rate $\eta$.
1:Choose $c\in(0,1]$, $\beta$ satisfying Eq. 9. (REC: start with $c=\frac{1}{2},\beta=1$).
2:for epoch $=1$ to $E$ do
@@
7:         Sample a mini-batch $B_{lb}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $((1-c)~{}\alpha_{y,a}+\frac{c}{A})\cdot B$
@@
12:     Update parameters $\theta$: $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{ext}$
@@ 4.2 GERNE for the Unknown Attributes Case
If the attributes are not known during training, it is not possible to directly sample biased and less biased batches. This is because the conditional attribute distribution, $p(a|y)$, is unknown. Hence, we follow the previous work [31, 28, 53] by training a standard ERM model $\tilde{f}$ and using its predictions to create pseudo-attributes $\tilde{a}$. By training $\tilde{f}$ using the biased batches, the model learns to rely on spurious correlations, leading to biased predictions. Leveraging these predictions, we can classify samples into hard examples (instances with low confidence) and easy examples (instances with high confidence). After training $\tilde{f}$, we select a threshold $t\in(0,1)$ and derive pseudo-attributes based on model predictions as follows: For each class $y$, we compute the predictions $\tilde{y}_{i}=p(y|x_{i})=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$. Then, we split the predictions into two non-empty subsets. The first subset contains the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ predictions, and the corresponding samples of these predictions represent group $\mathcal{X}_{y,\tilde{a}=1}$. The remaining samples forms the group $\mathcal{X}_{y,\tilde{a}=2}$. This process ensures that each set $\mathcal{X}_{y}$ is divided into two disjoint and non-empty groups. Consequently, the pseudo-attribute space consists of two values, denoted as $\mathcal{\tilde{A}}=\{1,2\}$ ($\tilde{A}=2$) with $\tilde{\mathcal{G}}=\mathcal{Y}\times\mathcal{\tilde{A}}$ replacing $\mathcal{G}$ in the unknown attribute case. $t$ is a hyperparameter and we outline the detailed steps for the unknown case in Algorithm 2 in Appendix D.
@@ 4.2.1 Tuning $\beta$ to control the unknown conditional distribution of an attribute $a$ in class $y$
After creating the pseudo-attributes and defining the pseudo-groups, suppose we form a new batch of size $B$ with $\gamma\cdot B$ samples uniformly sampled form group $(y,\tilde{a}=1)$ and $(1-\gamma)\cdot B$ samples from group $(y,\tilde{a}=2)$, for $\gamma\in[0,1]$. The conditioal attribute distribution of an attribute $a$ given $y$ in the new formed batch is given by:
$p_{B}(a|y)$ $=\sum_{\tilde{a}\in\mathcal{\tilde{A}}}p_{B}(\tilde{a}|y)\cdot p(a|\tilde{a},y)$ (12)
Because the max/min value of a linear program must occur at a vertex, we have:
$\forall\gamma\in[0,1],\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\leq p_{B}(a|y)\leq\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ (13)
This means that if: $\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)<\frac{1}{A}(\text{or }\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)>\frac{1}{A})$, then there is no value for $\gamma$ that allows creating a batch with $p_{B}(a|y)>\frac{1}{A}$ or $p_{B}(a|y)<\frac{1}{A})$ via sampling from the pseudo-groups.
Proposition 1. Creating a biased and less biased batch with the pseudo-attributes $\mathcal{\tilde{A}}=2$, and with $c,\beta$ as hyperparameter, we can simulate creating batches with more controllable conditional attribute distribution (i.e. $p_{B}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{B}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$). The proof is in Appendix E.
@@ 5 Experiments
The performance of our proposed approach, GERNE, is evaluated across five computer vision datasets and one natural language processing dataset, specifically: Colored MNIST (C-MNIST) [3, 27], Corrupted CIFAR-10 (C-CIFAR-10)[15, 31], Biased FFHQ (bFFHQ) [22, 27], Waterbird [45], CelebA [30], and CivilComments [5].
Based on the evaluation metric and experimental setup used, we divide these datasets into two groups: Datasets-1 and Datasets-2. Datasets-1 contains the first three datasets listed above, with the goal of assessing the performance of GERNE without any data augmentation methods. While Datasets-2 contains the remaining three datasets, with our implementation unified with [49]. In the following, we first describe our experimental setup and results for each group of datasets, and then we present an ablation study for hyperparameter tuning.
@@ 5.1 Experiments on Datasets-1
The C-MNIST dataset represents an extension of the MNIST dataset [26], incorporating colored digits. Each digit is highly correlated with a specific color, which constitutes its majority group. In the C-CIFAR-10 dataset, each category of images is corrupted with a specific type of texture noise [15]. The bFFHQ dataset comprises human face images, with ”age” and ”gender” as the target and spurious attributes, respectively. The majority of images depicting females are labeled as ”young” while the majority of images depicting males are labeled as ”old”.
@@
For both C-MNIST and C-CIFAR-10, we train with varying ratios of the minority to majority examples (0.5%, 1%, 2%, and 5%), and we follow the evaluation setup of [31, 29] using GBA on the test set as our evaluation metric. For bFFHQ, we train models with a 0.5% minority ratio and evaluate the performance based on the accuracy of the minority group in line with [27].
@@
We consider six baseline methods: For the known attribute case, we compare GERNE with Group DRO [39]. For the unknown attribute case, our baselines are ERM [44], JTT [28], LfF [31], DFA [27], and LC [29].
@@
We adopt the same model architectures as the baseline methods and utilize the SGD optimizer with a momentum of 0.9 and weight decay of 0.01 across all three datasets. Additional implementation details in Sec. F.1.
@@ Results.
Tab. 1 reports our results against baseline methods for the known and unknown attribute cases. The results of baseline methods are adopted from [29]. When the attributes are known, GERNE outperforms Group DRO by a significant margin on C-MNIST and C-CIFAR-10 datasets. The improvement in performance is ranging from about 5% on C-CIFAR-10 with 5% of minority group and up to 16% on C-MNIST with 1% of minority group. Furthermore, we show that we outperform the resampling method [16] ($c=1,\beta=0$) for all ratios. More discussion added to Appendix B. For bFFHQ, GERNE results in an improvement of over 7% in comparison with the resampling method. For the unknown attribute case, our method outperforms all the baseline methods except for C-MNIST with 5% of minority group where GERNE achieves the second-best accuracy, with only a 0.18% difference from the best result, while maintaining a lower standard deviation, despite not employing any data augmentation techniques.
Table 1: Performance comparison of GERNE and baseline methods on the C-MNIST, C-CIFAR-10, and bFFHQ datasets. We report the GBA (%) and standard deviation over three trials on the test set for C-MNIST and C-CIFAR-10, with varying ratios (%) of minority samples. For bFFHQ, we report the minority group accuracy (%). Baseline results are sourced from [29] as the same experimental settings are adopted. /$\checkmark$/$\times$ / × indicate the presence/absence of attribute information in the training set, respectively. The best results are marked in bold, and the second-best are underlined.
Methods Group Info C-MNIST C-CIFAR-10 bFFHQ 0.5 1 2 5 0.5 1 2 5 Group DRO 63.12 68.78 76.30 84.20 33.44 38.30 45.81 57.32 - GERNE (ours with $c=1,\beta=0$) 77.68 $\pm$0.89 84.36 $\pm$0.21 88.15 $\pm$0.11 91.98 $\pm$0.08 45.10 $\pm$0.60 50.08 $\pm$0.42 54.85 $\pm$0.30 62.16 $\pm$0.05 69.80 $\pm$1.33 GERNE (ours) 77.79 $\pm$0.90 84.47 $\pm$0.37 88.30 $\pm$0.20 92.16 $\pm$0.10 45.34 $\pm$0.60 50.84 $\pm$0.17 55.51 $\pm$0.10 62.40 $\pm$0.27 77.10 $\pm$0.90 ERM $\times$ 35.19 $\pm$3.49 52.09 $\pm$ 2.88 65.86 $\pm$ 3.59 82.17 $\pm$ 0.74 23.08 $\pm$ 1.25 28.52 $\pm$ 0.33 30.06 $\pm$ 0.71 39.42 $\pm$ 0.64 56.70 $\pm$ 2.70 JTT $\times$ 53.03 $\pm$ 3.89 62.90 $\pm$ 3.01 74.23 $\pm$ 3.21 84.03 $\pm$ 1.10 24.73 $\pm$ 0.60 26.90 $\pm$ 0.31 33.40 $\pm$ 1.06 42.20 $\pm$ 0.31 65.30 $\pm$ 2.50 LfF $\times$ 52.50 $\pm$ 2.43 61.89 $\pm$ 4.97 71.03 $\pm$ 1.14 84.79 $\pm$ 1.09 28.57 $\pm$ 1.30 33.07 $\pm$ 0.77 39.91 $\pm$ 1.30 50.27 $\pm$ 1.56 62.20 $\pm$ 1.60 DFA $\times$ 65.22 $\pm$ 4.41 81.73 $\pm$ 2.34 84.79 $\pm$ 0.95 89.66 $\pm$ 1.09 29.75 $\pm$ 0.71 36.49 $\pm$ 1.79 41.78 $\pm$ 2.29 51.13 $\pm$ 1.28 63.90 $\pm$ 0.30 LC $\times$ 71.25 $\pm$ 3.17 82.25 $\pm$ 2.11 86.21 $\pm$ 1.02 91.16 $\pm$ 0.97 34.56 $\pm$ 0.69 37.34 $\pm$ 1.26 47.81 $\pm$ 2.00 54.55 $\pm$ 1.26 69.67 $\pm$ 1.40 GERNE (ours) $\times$ 77.25 $\pm$ 0.17 83.98 $\pm$ 0.26 87.41 $\pm$ 0.31 90.98 $\pm$ 0.13 39.90$\pm$ 0.48 45.60$\pm$ 0.23 50.19$\pm$ 0.18 56.53 $\pm$ 0.32 70.87 $\pm$ 0.61

### v2

@@ 4.1.3 GERNE as a general framework for debiasing
We provide the full proof in Appendix A.
Based on Eq. 8, we can establish the link between GERNE and other methods for different values of $\beta,c$:
    For $\beta=-1$, $\mathcal{L}_{ext}=\mathcal{L}_{b}$ and GERNE is equivalent to class-balanced ERM method.
    For $c=1$ and $\beta=0$, $p_{ext}(a|y)=\frac{1}{A}$, and GERNE matches Resampling  [16], which samples equally from all groups ($B_{rs}$ in Fig. 1(b), with gradient of the loss computed on it denoted as $\nabla_{\theta}\mathcal{L}_{rs}$ in Fig. 1(c)).
    For $c\cdot(\beta+1)=1$, we also have $p_{ext}(a|y)=\frac{1}{A}$, and $\mathcal{L}_{ext}$ is, in expectation, equivalent to $\mathcal{L}_{rs}$. However, their loss variances differ. In fact, GERNE permits controlling the variance of its loss through its hyperparameters ($c,\beta$), which may help escape sharp minima  [1] and improve generalization  [23]. The derivation of the variance of GERNE’s loss is detailed in Appendix B.
    For $c\cdot(\beta+1)>1$, $p_{ext}(a|y)>\frac{1}{A}\text{ if }\alpha_{ya}<\frac{1}{A}$ (also $p_{ext}(a|y)<\frac{1}{A}\text{ if }\alpha_{ya}>\frac{1}{A}$). In this case, GERNE simulates batches where the underrepresented groups (i.e., those with $\alpha_{ya}<\frac{1}{A}$) are oversampled.
@@ 4.1.4 Upper and lower bounds of $\beta$
Having $p_{ext}(a|y)$ in Eq. 8 within $[0,1]$, $\beta$ should satisfy:
$\begin{split}\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\min(i_{ya}^{1},i_{ya}^{2})\leq\beta\leq\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2}),\end{split}$ (9)
where: $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$.
These bounds are used when tuning $\beta$. In Appendix C, we simplify these bounds to $\beta_{\min},\beta_{\max}]=[-1,i_{y^{\prime\prime}a^{\prime\prime}}^{1}]$, where $(y^{\prime\prime},a^{\prime\prime})=\arg\max_{(y,a)\in\mathcal{G}}\alpha_{ya}$. Note that $\beta$ doesn’t affect $p_{ext}(a|y)$ for $\alpha_{ya}=\frac{1}{A}$ according to Eq. 8.
4.1.5 Tuning $\beta$ to minimize the risk for worst-case group
Eq. 5 can be rewritten as follows (detailed in Appendix A):
$\mathcal{L}_{ext}=\frac{1}{K}\cdot\sum_{g=(y,a)\in\mathcal{G}}p_{ext}(a|y)(\beta)\cdot L_{g},$ (10)
@@ 4.1.5 Tuning $\beta$ to minimize the risk for worst-case group
In the presence of spurious correlations, minority or less-represented groups often experience higher risks, primarily due to the model’s limited exposure to these groups during training  [20]. Taking this into consideration, we define $g^{\prime}=(y^{\prime},a^{\prime})=\arg\min_{(y,a)\in\mathcal{G}}\alpha_{ya}$. Since $L_{g^{\prime}}$ is weighted by $p_{ext}(a^{\prime}|y^{\prime})$, increasing $\beta$ beyond $\frac{1}{c}-1$ assigns more weight to $L_{g^{\prime}}$ in Eq. 10 than any other group loss (all groups’ losses are equally weighted when $c\cdot(\beta+1)=1$). This increase in $\beta$ encourages the model to prioritize reducing the loss of the underrepresented group $g^{\prime}$ during training, therefore minimizing the risk for the worst-case group.
We outline the detailed steps of our approach for the known attribute case in Algorithm 1.
@@
Input: $\mathcal{X}_{y,a}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$ and $a\in\mathcal{A}$, $f$ with initial $\theta=\theta_{0}$, #\# epochs $E$, batch size per label $B$, #\# classes $K$, #\# attributes $A$, learning rate $\eta$.
1:Choose $c\in(0,1]$ and $\beta\in[\beta_{\min},\beta_{\max}]$ via grid search.
2:for epoch =1=1 to $E$ do
@@
7:   Sample a mini-batch $B_{lb}^{y,a}=\{(x,y)\}\subseteq\mathcal{X}_{y,a}$ of size $((1-c)\cdot\alpha_{y,a}+\frac{c}{A})\cdot B$
@@
12:  Update parameters (SGD): $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}_{ext}$
@@ 4.2 GERNE for the Unknown Attributes Case
If the attributes are unavailable during training, it is not possible to directly sample less biased batches. To address this, we follow the previous work  [31, 28, 54] by training a standard ERM model $\tilde{f}$ and using its predictions to create pseudo-attributes $\tilde{a}$. Since $\tilde{f}$ is trained on biased batches, it tends to rely on spurious correlations, resulting in biased predictions. Leveraging these predictions, we classify samples into easy—those with high-confidence predictions, where the spurious correlations likely hold— and hard — those with low-confidence predictions, where the spurious correlations may not hold. After training $\tilde{f}$, we select a threshold $t\in(0,1)$ and construct pseudo-attributes based on model predictions as follows: For each class $y$, we compute the predictions $\tilde{y}_{i}=p(y|x_{i})=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$. We then split them into two non-empty subsets: The first subset contains the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ values, and the corresponding samples form the group $\mathcal{X}_{y,\tilde{a}=1}$. The remaining samples forms the group $\mathcal{X}_{y,\tilde{a}=2}$. This process ensures that each set $\mathcal{X}_{y}$ is divided into two disjoint and non-empty groups. Consequently, the pseudo-attribute space consists of two values, denoted as $\mathcal{\tilde{A}}=\{1,2\}$ (i.e., $\tilde{A}=2$) with $\tilde{\mathcal{G}}=\mathcal{Y}\times\mathcal{\tilde{A}}$ replacing $\mathcal{G}$ in the unknown attribute case. $t$ is a hyperparameter, and we outline the detailed steps of GERNE for the unknown case in Appendix D.
@@ 4.2.1 Tuning $\beta$ to control the unknown conditional distribution of an attribute $a$ in class $y$
After creating the pseudo-attributes and defining the pseudo-groups, we consider forming a new batch of size $B$ by uniformly sampling $\gamma\cdot B$ examples from group $\mathcal{X}_{y,\tilde{a}=1}$ and $(1-\gamma)\cdot B$ examples from group $\mathcal{X}_{y,\tilde{a}=2}$, where $\gamma\in[0,1],\gamma\cdot B\in\mathbb{N}$. The resulting conditional distribution of an attribute $a$ given $y$ in the constructed batch is:
$p_{B}(a|y)$ $=\sum_{\tilde{a}\in\mathcal{\tilde{A}}}p_{B}(\tilde{a}|y)\cdot p(a|\tilde{a},y).$ (12)
Because the max/min value of a linear program must occur at a vertex, we have for $p(a|\tilde{a},y)=p_{\tilde{a},y}(a)$:
$\forall\gamma\in[0,1],\min_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)\leq p_{B}(a|y)\leq\max_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a).$ (13)
This means that if: $\max_{\tilde{a}}p_{\tilde{a},y}(a)<\frac{1}{A}$$(\min_{\tilde{a}}p_{\tilde{a},y}(a)>\frac{1}{A})$, then there is no value for $\gamma$ can yield a batch with $p_{B}(a|y)>\frac{1}{A}$$(p_{B}(a|y)<\frac{1}{A})$ via sampling from the pseudo-groups.
Proposition 1.
In case of unknown attributes, GERNE can simulate creating batches with more controllable conditional attribute distribution (i.e., $p_{B}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)\text{ or }p_{B}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p_{\tilde{a},y}(a)$). We provide the proof of this proposition in Appendix E.
@@ 5 Experiments
To evaluate the general applicability of GERNE, we assess its performance across five computer vision and one natural language processing benchmarks: Colored MNIST (C-MNIST)  [3, 27], Corrupted CIFAR-10 (C-CIFAR-10) [15, 31], Biased FFHQ (bFFHQ)  [22, 27], Waterbird  [46], CelebA  [30], and CivilComments  [5]. We categorize these datasets into two groups: Datasets-1 and Datasets-2. Datasets-1 comprises the first three datasets mentioned above and is used to evaluate GERNE’s performance without data augmentation. Datasets-2 consists of the remaining three datasets, for which we follow the experimental setup described in  [50] to ensure a fair comparison.
@@ 5.1 Experiments on Datasets-1
C-MNIST is an extension of the MNIST dataset  [26] where each digit class is predominantly associated with a specific color. This introduces a spurious correlation between the digit label (target) and color (attribute). C-CIFAR-10 modifies CIFAR-10 by applying specific texture patterns to each object class  [15], making texture a spurious feature. Both C-MNIST and C-CIFAR-10 include versions with varying degrees of spurious correlation, reflected by the minority group ratios of 0.5%, 1%, 2%, and 5% in the training and validation sets. The bFFHQ dataset comprises human face images, with “age” and “gender” as the target and spurious attributes, respectively. The majority of female faces are young, while the majority of males are old. The minority group ratio in the training set is 0.5%.
@@
We follow the evaluation protocols of prior work [31, 29, 27]. For C-MNIST and C-CIFAR-10, we report Group-Balanced Accuracy (GBA) on the test set. For bFFHQ, we evaluate performance based on the accuracy of the minority group.
@@
For the known attribute case, we compare GERNE with Group DRO  [39] and Resampling  [16]. For the unknown attribute case, we consider ERM  [45], JTT  [28], LfF  [31], DFA  [27], LC  [29], and DeNetDM  [44].
@@
We adopt the same model architectures as the baselines and use SGD optimizer across all three datasets. More details are provided in Sec. F.1.
Table 1: Performance comparison of GERNE and baselines on the C-MNIST, C-CIFAR-10, and bFFHQ datasets. We report GBA (%) with standard deviation over three trials for C-MNIST and C-CIFAR-10 across varying minority ratios, and minority group accuracy (%) for bFFHQ. DeNetDM results are from [44], and Resampling results are generated using GERNE with $c=1,\beta=0$. All other results are from  [29]. $\checkmark/\times$ indicates known/unknown training attributes. The best results are marked in bold, and the second-best are underlined.
[…]
@@ Results.
Tab. 1 compares GERNE with baselines for both known and unknown attribute cases. All baseline results are adopted from  [29], except DeNetDM, which is sourced from  [44]. When the attributes are known, GERNE outperforms Group DRO by a significant margin on C-MNIST and C-CIFAR-10 datasets. The improvement in performance ranges from about 5% on C-CIFAR-10 with 5% of minority group and up to 16% on C-MNIST with 1% of minority group. Furthermore, GERNE outperforms Resampling  [16] by over 13% on bFFHQ and consistently surpasses it across all other ratios. Our explanation behind GERNE’s superior performance over Resampling is that the latter tends to present the majority and minority groups equally in the sampled batches during training, and the model $f$ tends to prioritize learning the easy-to-learn spurious features associated with the majority group (e.g., the colors in C-MNIST), leading to learning biased representation and poorer generalization. In contrast, GERNE undermines learning the spurious features by directing the learning process more in the debiasing direction, thanks to the extrapolation factor. For the unknown attribute case, GERNE outperforms all baselines, except on C‑MNIST with 5% of minority group (ranks second), while maintaining a lower standard deviation. At this 5% minority ratio, LC achieves slightly higher accuracy—likely benefiting from its use of data augmentation to increase the diversity of the samples in the minority group. We exclude DeNetDM’s results on C-MNIST, as the authors use a different version of this dataset.